swinn:

Which of the following is not a true statement? these are my options a=v2-v1/t Ac=v2/r t=Δa/v Δv = at

Warriorz13:

These links take me to nothing

XioGonz:

They are blank

darkknight:

So acceleration is the change in velocity over time, so the first option is correct. so a = b/t, we can rearrange so that at = v, option d is correct. Solvng for t we have t = v/a, not t = a/v, This is incorrect, So t=Δa/v is the incorrect one and it is your answer.